OpenStudy (anonymous):

need help with ln e^4

OpenStudy (anonymous):

\[x=\ln e ^{4}\]\[e ^{x}=e ^{4}\]\[x =4\]

OpenStudy (anonymous):

how do they get from \[x =\ln e ^{4}\] to \[e ^{x}=e ^{4}\] sorry if its dumb question. I am aware that \[\ln e ^{4}=\log_{e}e ^{4} =4\] I am just wondering about the algebra used here.

OpenStudy (anonymous):

they could have skipped the second step and immediately get x=4. but what they did was they took exponential of both sides which resulted in that second line

OpenStudy (anonymous):

what do u mean exponential of both sides

OpenStudy (anonymous):

wait.. cause x = ln x ?

OpenStudy (anonymous):

no. \[x=\ln \exp(4)\] \[\exp(x)=\exp \ln \exp(4)\] so \[\exp(x)=\exp(x)\]

OpenStudy (anonymous):

**\[\exp(x)=\exp(4)\]

OpenStudy (anonymous):

\[e ^{x}=e ^{\ln e ^{4}}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

what is the rule that allows you to drop the ln from the right side exponent?

OpenStudy (anonymous):

exp and ln are inverse functions of each other so exp and ln cancel each other

OpenStudy (anonymous):

wooo awesome.. thats what i am missing in my head. so u dont see ln in exponentials ;p

OpenStudy (anonymous):

thanks for sticking with me.. my brain gets so stuck on easy stuff... i just feel like i need to know every possible way